OpenStudy (anonymous):

logarithms with integration. does the ln(-3) cancel?

OpenStudy (anonymous):

\[\frac{ 1 }{ 6 }(\ln 5-\ln11-\ln \left| -3 \right|+\ln3)\]

OpenStudy (anonymous):

how does that become \[\frac{ 1 }{ 6 } \ln \frac{ 15 }{ 11 }\]? does the ln -3 cancel?

OpenStudy (anonymous):

i kept getting 15/33...so im assuming the -3 canceled?

OpenStudy (anonymous):

i know that... but doesnt the absolute value make it positive?

OpenStudy (callisto):

\[\frac{ 1 }{ 6 }(\ln 5-\ln11-\ln \left| -3 \right|+\ln3) \]\[=\frac{ 1 }{ 6 }(\ln 5-\ln11-\ln 3+\ln3)\]\[=\frac{ 1 }{ 6 }(\ln 5-\ln11 +(\ln3 -\ln3))\]\[=\frac{ 1 }{ 6 }(\ln 5-\ln11 +0)\]\[=\frac{ 1 }{ 6 }(\ln 5-\ln11)\] This is what I can get :\

OpenStudy (callisto):

\[=\frac{ 1 }{ 6 }\ln \frac{5}{11}\]

OpenStudy (anonymous):

ooooh! i thought the - ln3 would go to the denominator

OpenStudy (callisto):

Problem solved?

OpenStudy (anonymous):

yes thanks a bunch

OpenStudy (callisto):

Welcome :) Sorry for reading your question wrong earlier!

OpenStudy (anonymous):

no worries